OpenStudy (anonymous):

find the domain of validity for each identity sin0(imagine the 0s has a line through it)+cos0/sin 0=1+cos0 Please explain I am totally lost on this I know that it is true but i don't know anything else is that all I need to know I am really confused please help!

OpenStudy (lunymoony):

I have no idea, sorry. Where's a math/trigonometry genius when you need them!

OpenStudy (anonymous):

I am really confused on this

OpenStudy (lunymoony):

I did these once, but I sucked at them and I have no idea how I got through them

OpenStudy (loser66):

\(sin(\theta)+\dfrac{cos(\theta)}{sin(\theta)}= 1+ cos (\theta)\), right?

OpenStudy (loser66):

and we have to prove it, right?

OpenStudy (anonymous):

yep

OpenStudy (loser66):

It is not identity, how to prove?? counterexample, \(\theta = \pi/6\)

OpenStudy (loser66):

Left hand side: \(1/2+ \sqrt3=\dfrac{1+2\sqrt3}{2}\) Righ hand side :\(1+\sqrt3/2 =\dfrac{2+\sqrt3}{2}\)

OpenStudy (loser66):

Hence \(LHS\neq RHS\)

OpenStudy (anonymous):

i am going to make a guess `

OpenStudy (anonymous):

no, i am not

OpenStudy (anonymous):

but i am almost certain there is a typo in the question